# Monocular Person Localization under Camera Ego-motion

### v1

@@ IV-A Baselines
To evaluate the localization accuracy of the proposed method, we conduct a comparison with geo-model-based and deep-learning-based methods introduced in Sec. II-A and Sec. II-C respectively. The baselines are named as follows: Geo-model-based:
Koide’s Method[3] locates the person by the neck observation under the assumption of a fixed camera.
Ye’s Method[4] extends Koide’s method to use observations of four points separately to handle occlusions, still restricted to a fixed camera.
@@
Depth Anything[21] outputs relative depth map of the scene. We use the ground truth depth of the person in the first scene as the metric, and compute the average depth of all the joints as the person’s distance from the camera.
Multi-HMR[25] focuses on human mesh recovery but also estimates the absolute distance between the pelvis (the center of human hips) and the camera through a neural network.
@@ IV-B Datasets
FieldSAFE[43], which is the closest to our scenario to the best of our knowledge. This dataset consists of approximately two hours of multi-sensor data collected from a tractor-mounted system in a grass-mowing scenario and provides precise localization of the observed moving pedestrians.
KITTI[35] where most deep-learning-based person localization models are trained on. However, it should be noted that the KITTI dataset, which covers autonomous driving scenarios, does not include data distributions with continuous and severe camera ego-motion from a robot’s view.
To evaluate the proposed method and to support further research on this topic, we exhaustively build a dataset and make it public.
The dataset is recorded from the quadruped robot Unitree Go1[7], illustrated in Fig. 4(a) (refer to Sec. IV-C). It consists of three scenarios: Turning Head, Indoor Slope and Rugged Lawn. Fig. 4(b), 4(c), 4(d) are snapshots of them respectively. In the latter two scenarios, we run our RPF system in real time while collecting data from pin-hole, fisheye and panoramic cameras. In the Rugged Lawn scenario, the Go1 robot followed the target person for six minutes over a distance of 354 meters without any pauses. We utilized UWB to measure the person’s distance from the camera as the ground truth. In the first two scenarios, we used a motion capture system to record the relative poses between the person and the robot as ground truth.
Our dataset includes data collected at close range, with large-angle ego-motion and continuous observations of a real walking person from a low viewing angle. Table I illustrates the range of the data distribution compared to other datasets.
| Dataset | KITTI[35] | FieldSAFE[43] | Our Dataset |
@@ IV-C Platform and Implementation Details
We use quadruped robot Unitree Go1[7] as our experimental platform (shown in Fig. 4(a)) which equips with: an onboard mini PC Intel NUC 11 containing CPU i7-1165g7 and GPU RTX 2060-mobile (6G), a mounted UWB distance sensor with 0.1m accuracy (Nooploop LinkTrack AOA) for ground truth distance measurement, and two forward-facing cameras (Intel RealSense D435i and T265), capturing 640*480 pin-hole and 848*800 fisheye grayscale images with global shutter at 30 Hz, respectively.
@@
It is worth mentioning that, compared to the quadruped platforms used in existing RPF methods[8, 10, 11], our Go1 robot is smaller in size and has a higher step frequency. As a result, it performs fierce ego-motion in rough terrains.
We evaluate all the methods except for Multi-HMR[25] on the onboard mini PC of Go1 robot. We adopt the largest model for Depth Anything[21] and Multi-HMR[25]. For methods that use 2D human joint detection as input, the measurements are uniformly obtained through the method specified in Sec. III-A. We leverage NVIDIA’s TensorRT technology to accelerate the inference of YOLOX[36] and AlphaPose[37] for real-time performance. As for Multi-HMR, due to the massive model parameter size, the mini PC is unable to run it; instead, we conduct the testing offline on a separate PC equipped with RTX 3070 (8G) and an Intel i7-12700F.
@@ IV-D Evaluation and Results
For localization accuracy, we uniformly adopt the pelvis (the center of human hips) in the camera frame to represent the 3D location of a human. Additionally, we calculate the distance of the human relative to the optical center, which is equal to the L2 norm of the location. We employ the Average Location Error (ALE) and Average Distance Error (ADE) to evaluate the accuracy of localization as [23, 4] did. Moreover, for FieldSAFE[43] and Rugged Lawn scenario in our dataset with continuous camera ego-motion, we introduce the Variance of Location Error (VLE) and Variance of Distance Error (VDE) as additional metrics to assess the stability of localization. Since all the baselines only support pin-hole camera model, we use the pin-hole images in our dataset for evaluation across baselines. Fig. 5(a), 5(b) show screenshots of our method running on public datasets. Fig. 5(c) and Fig. 1 show our method running on equirectangular and fisheye images, respectively. Fig. 5(d) shows our method still works under partial observation (discussed in Sec. IV-E).
@@
Additionally, a box plot illustrating the distance error of our method compared to baselines in the Rugged Lawn scenario is presented in Fig. 6. Our method demonstrates more accurate and stable distance estimation throughout the six-minute RPF process, even under continuous and fierce camera ego-motion. Specifically, we extract a clip of the Rugged Lawn scene (from 10s to 57s) and plot the person distance estimation curves for both deep-learning-based (Fig. 7(a)) and geo-model-based methods (Fig. 7(b)).
| Turning Head | Indoor Slope | Rugged Lawn | FieldSAFE[43] | KITTI[35] | |
@@
| Koide’s Method[3] | 0.396 | 0.289 | 0.3 / 0.3 | 1.924 / 5.012 | 1.451 |
| Ye’s Method[4] | 0.294 | 0.261 | 0.3 / 0.3 | 1.856 / 3.952 | 1.420 |
| MonoLoco++[23] | 0.820 | 0.510 | 0.6 / 0.2 | 4.152 / 4.705 | 0.940 |
| Depth Anything[21] | 0.571 | 0.523 | 0.5 / 0.6 | 1.528 / 1.022 | 2.963 |
| Multi-HMR[25] | 0.493 | 0.254 | 0.4 / 0.3 | 3.066 / 0.424 | 1.520 |
@@
As shown in Fig. 7(a), due to weak generalization, the deep-learning-based methods struggle to accurately reflect the trend of person distance changes during continuous RPF. This is particularly evident with Monoloco++[23], which was trained on the KITTI dataset[35] that lacks data distribution for a person below three meters from the camera, making it difficult to predict accurate distances of a close person in Rugged Lawn. From Fig. 7(b), the geo-model-based methods generally capture the trend of distance changes. However, when the camera undergoes large-angle ego-motion, methods except ours have significant errors, which could pose safety concerns for an RPF system that relies on localization for robot speed control. This further highlights the robustness and necessity of our method.
@@ IV-E Discussion
Partial observation. Our method still works even when only three points in are observed, demonstrating strong robustness against partial observations of a human body, such as the scenario in Fig. 5(d). As shown in Table II, when the neck’s or ankle’s observation is removed across datasets, our method maintains high localization accuracy.
Performance against posture change. As shown in Fig. 7(b), the distance estimated by our method consistently oscillates around the ground truth at a certain frequency. This oscillation arises because the human body undergoes periodic deformation during walking. Such fluctuations in distance can be easily smoothed out by a filter. However, when the human breaks the upright assumption such as sitting down, our method fails as most geo-model-based methods do. This issue can be engineerly resolved by incorporating a human posture recognition module that filters out non-upright postures.
@@ V CONCLUSIONS
In this paper, we propose an optimization-based method to locate a target person under camera ego-motion. By representing an upright human with a four-point model, our method jointly estimates camera attitude and person location. A Robot Person Following (RPF) system is proposed based on our method that achieves real-time (40 FPS on an onboard mini PC) monocular person localization and tracking from an agile quadruped’s view in rough terrains. We create and make public a dataset recorded from the quadruped to supplement the research of monocular person localization under camera ego-motion, which is critical for human-robot interaction such as RPF[1] and egocentric surveillance systems[2]. Experiments on both public datasets and our dataset demonstrate the effectiveness of our method compared to deep-learning-based and geo-model-based baselines. Future work will involve utilizing more sophisticated models to represent the human body, thereby enabling monocular person localization across various human postures. Additionally, integrating the estimation of the ground plane where a person is standing may enhance the accuracy of person localization in 3D space.

### v2

@@ IV-A Baselines
To evaluate the localization accuracy of the proposed method, we conduct a comparison with geo-model-based and deep-learning-based methods introduced in Sec. II-A and Sec. II-C, respectively. The baselines are named as follows: Geo-model-based:
Koide’s Method[koide2020monocular] locates the person by the neck observation under the assumption of a fixed camera.
Ye’s Method[ye2023robot] extends Koide’s method to use observations of four points separately to handle occlusions, still restricted to a fixed camera.
@@
MonoLoco++[monoloco++] estimates the 3D coordinates of the bounding box’s center of the person in camera frame by a neural network trained on the KITTI dataset[kitti].
Depth Anything[depth_anything] outputs relative depth map of the scene. We use the ground truth depth of the person in the first scene as the metric, and compute the average depth of all the joints as the person’s distance from the camera.
Multi-HMR[multi-hmr2024] focuses on human mesh recovery but also estimates the absolute distance between the pelvis (the center of human hips) and the camera through a neural network.
@@ IV-B Datasets
FieldSAFE[fieldsafe], which is the closest to our scenario to the best of our knowledge. This dataset consists of approximately two hours of multi-sensor data collected from a tractor-mounted system in a grass-mowing scenario and provides precise localization of the observed moving pedestrians.
KITTI[kitti] where most deep-learning-based person localization models are trained on. However, it should be noted that the KITTI dataset, which covers autonomous driving scenarios, does not include data distributions with severe camera ego-motion from a robot’s view.
To evaluate the proposed method and to support further research on this topic, we constructed and released a dataset named RPF-Quadruped.
RPF-Quadruped dataset is recorded from the quadruped robot Unitree Go1[unitree], illustrated in Fig. 4(a) (refer to Sec. IV-C). It consists of three scenarios: Turning Head, Indoor Slope, and Rugged Lawn. Fig. 4(b), 4(c), 4(d) are snapshots of them, respectively. In the latter two scenarios, we run our RPF system in real time while collecting data from pin-hole, fisheye, and panoramic cameras. In the Rugged Lawn scenario, the Go1 robot followed the target person for six minutes over a distance of 354 meters without any pauses. We utilized UWB to measure the person’s distance from the camera as the ground truth. In the first two scenarios, we used a motion capture system to record the relative poses between the person and the robot as ground truth.
RPF-Quadruped dataset includes data collected at close range, with large-angle ego-motion and continuous observations of a real walking person from a low viewing angle. Table I illustrates the range of the data distribution compared to other datasets.
| Dataset | KITTI[kitti] | FieldSAFE[fieldsafe] | RPF-Quadruped |
@@ IV-C Platform and Implementation Details
We use a quadruped robot Unitree Go1[unitree] as our experimental platform (shown in Fig. 4(a)), which is equipped with: an onboard mini PC Intel NUC 11 containing CPU i7-1165g7 and GPU RTX 2060-mobile (6G), a mounted UWB distance sensor with 0.1m accuracy (Nooploop LinkTrack AOA) for ground truth distance measurement, and two forward-facing cameras (Intel RealSense D435i and T265), capturing 640*480 pin-hole and 848*800 fisheye grayscale images with global shutter at 30 Hz, respectively.
@@
It is worth mentioning that, compared to the quadruped platforms used in existing RPF methods[zhang2021efficient, Multi-modal, 3D-pose-korean], our Go1 robot is smaller in size and has a higher step frequency. As a result, it performs severe ego-motion in rough terrain.
We evaluate all the methods except for Multi-HMR[multi-hmr2024] on the onboard mini PC of the Go1 robot. We adopt the largest model for Depth Anything[depth_anything] and Multi-HMR[multi-hmr2024]. For methods that use 2D human joint detection as input, the measurements are uniformly obtained through the method specified in Sec. III-A. We leverage NVIDIA’s TensorRT technology to accelerate the inference of YOLOX[yolox] and AlphaPose[fang2022alphapose] for real-time performance. As for Multi-HMR, due to the massive model parameter size, the mini PC is unable to run it; instead, we conduct the testing offline on a separate PC equipped with RTX 3070 (8G) and an Intel i7-12700F.
@@ IV-D Evaluation and Results
For localization accuracy, we uniformly adopt the pelvis (the center of human hips) in the camera frame to represent the 3D location of a human. Additionally, we calculate the distance of the human relative to the optical center, which is equal to the L2 norm of the location. We employ the Average Location Error (ALE) and Average Distance Error (ADE) to evaluate the accuracy of localization as [monoloco++, ye2023robot] did. Moreover, for FieldSAFE[fieldsafe] and Rugged Lawn scenario in our dataset with continuous camera ego-motion, we introduce the Variance of Location Error (VLE) and Variance of Distance Error (VDE) as additional metrics to assess the stability of localization. Since all the baselines only support the pin-hole camera model, we use the pin-hole images in our dataset for evaluation across baselines. Fig. 5(a), 5(b) show screenshots of our method running on public datasets. Fig. 5(c) and Fig. 1 show our method running on equirectangular and fisheye images, respectively. Fig. 5(d) shows our method still works under partial observation (discussed in Sec. IV-E).
@@
Additionally, a box plot illustrating the distance error of our method compared to baselines in the Rugged Lawn scenario is presented in Fig. 6. Our method demonstrates more accurate and stable distance estimation throughout the six-minute RPF process, even under continuous and severe camera ego-motion. Specifically, we extract a clip of the Rugged Lawn scene (from 10s to 57s) and plot the person distance estimation curves for both deep-learning-based (Fig. 7(a)) and geo-model-based methods (Fig. 7(b)).
| Turning Head | Indoor Slope | Rugged Lawn | FieldSAFE[fieldsafe] | KITTI[kitti] | |
@@
| Koide’s Method[koide2020monocular] | 0.396 | 0.289 | 0.3 / 0.3 | 1.924 / 5.012 | 1.451 |
| Ye’s Method[ye2023robot] | 0.294 | 0.261 | 0.3 / 0.3 | 1.856 / 3.952 | 1.420 |
| MonoLoco++[monoloco++] | 0.820 | 0.510 | 0.6 / 0.2 | 4.152 / 4.705 | 0.940 |
| Depth Anything[depth_anything] | 0.571 | 0.523 | 0.5 / 0.6 | 1.528 / 1.022 | 2.963 |
| Multi-HMR[multi-hmr2024] | 0.493 | 0.254 | 0.4 / 0.3 | 3.066 / 0.424 | 1.520 |
@@
As shown in Fig. 7(a), due to weak generalization, the deep-learning-based methods struggle to accurately reflect the trend of person distance changes during continuous RPF. This is particularly evident with Monoloco++[monoloco++], which was trained on the KITTI dataset[kitti] that lacks data distribution for a person below three meters from the camera, making it difficult to predict accurate distances of a close person in Rugged Lawn. From Fig. 7(b), the geo-model-based methods generally capture the trend of distance changes. However, when the camera undergoes large-angle ego-motion, methods except ours have significant errors, which could pose safety concerns for an RPF system that relies on localization for robot speed control. This further highlights the robustness and necessity of our method.
| Method | Preprocessing (s) | Estimation (s) | Total (s) |
| Koide’s Method[koide2020monocular] | 0.02 | 0.0006 | 0.0206 |
| Ye’s Method[ye2023robot] | 0.02 | 0.0008 | 0.0208 |
| MonoLoco++[monoloco++] | 0.02 | 0.09 | 0.11 |
| Depth Anything[depth_anything] | / | 0.23 | 0.23 |
| Multi-HMR[multi-hmr2024]∗ | / | 1.24 | 1.24 |
| Ours | 0.02 | 0.005 | 0.025 |
@@ IV-E Discussion
Performance against posture change. As shown in Fig. 7(b), the distance estimated by our method consistently oscillates around the ground truth at a certain frequency. This oscillation arises because the human body undergoes periodic deformation during walking. Such fluctuations in distance can be easily smoothed out by a filter. However, when the human breaks the upright assumption, such as sitting down, our method fails as most geo-model-based methods do. This issue can be engineerly resolved by incorporating a human posture recognition module that filters out non-upright postures.
@@ V CONCLUSIONS
In this paper, we propose an optimization-based method to locate a target person under camera ego-motion. By representing an upright human with a four-point model, our method jointly estimates camera attitude and person location. A Robot Person Following (RPF) system is proposed based on our method that achieves real-time (40 FPS on an onboard mini PC) monocular person localization and tracking from an agile quadruped’s view in rough terrains. We create and make public a dataset recorded from the quadruped to supplement the research of monocular person localization under camera ego-motion, which is critical for human-robot interaction such as RPF[islam2019person] and egocentric surveillance systems[anomaly]. Experiments on both public datasets and our dataset demonstrate the effectiveness of our method compared to deep-learning-based and geo-model-based baselines. Future work will involve extending our method by employing expressive human models to handle a wider range of postures. The localization accuracy could also be improved by explicitly estimating the ground plane on which the person stands. Extensive evaluations can be conducted on large-scale egocentric datasets with dense crowds, such as TPT-bench[tpt-bench].